OpenStudy (alli6106):

When its digits are reversed, a positive two-digit integer is increased by 20%. What is the integer?

OpenStudy (amistre64):

let the number be: 10a + b (10b+a)/(10a+b) = 1.20

OpenStudy (amistre64):

or 10b+a = (10a+b)(.20) 10b+a = 2a+b(.20) 10b = a+b(.20) 9.8b = a 9.8 = a/b

OpenStudy (amistre64):

bah, 1.20 was more better

OpenStudy (alli6106):

what is the integer

OpenStudy (amistre64):

work the process, im not here to do your work and thinking for you am i?

OpenStudy (alli6106):

ok

OpenStudy (amistre64):

yeah, it works fine 10b + a = (10a + b)(1.20) work it out into a/b and youll get the results